# DenoiseReID: Denoising Model for Representation Learning of Person Re-Identification

### v1

@@ Appendix D Limitations
Our proposed method FEFDFA improves the accuracy of current mainstream backbones while ensuring label-free and no additional computational costs, and it has been experimentally verified to be generalizable in multiple image tasks. However, from the experimental results, it can be found that our method has limited improvement in model accuracy when generalized to general tasks, and in order to fuse the parameters of the denoising layer and the feature extraction layer, only one or two steps of denoise for each denoising layer, and the number of denoising layers is not more than that of the feature extraction layer, which limits the denoising intensity. We will continue to explore how to further improve the accuracy of the model without adding additional inference time costs or only adding a small number of additional parameters.

### v2

@@ Appendix F Limitations
Our proposed method DenoiseRep improves the accuracy of current mainstream backbones while ensuring label-free and no additional computational costs, and it has been experimentally verified to be generalizable in multiple image tasks. However, from the experimental results, it can be found that our method has limited improvement in model accuracy when generalized to general tasks, and in order to fuse the parameters of the denoising layer and the feature extraction layer, only one or two steps of denoise for each denoising layer, and the number of denoising layers is not more than that of the feature extraction layer, which limits the denoising intensity. We will continue to explore how to further improve the accuracy of the model without adding additional inference time costs or only adding a small number of additional parameters.